OpenStudy (anonymous):

CAN SOMEONE PLEASE PLEASE HELP??!! I've been asking for the past 3 days... Does anyone get this? It keeps saying my answer is wrong., ndicate the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(1, 3) and B(0, 2)

OpenStudy (anonymous):

 was my answer

OpenStudy (anonymous):

What did you get as the slope?

OpenStudy (anonymous):

find the slope, then put it in the equation y-y1=mx-x1

OpenStudy (anonymous):

my slope was 1

OpenStudy (anonymous):

slope=y2-y1/x2-x1

OpenStudy (anonymous):